OpenStudy (anonymous):

Can someone plz help me get started on these fractions f(x)=3x-8 div 4x+20

OpenStudy (anonymous):

What do you need?

OpenStudy (saifoo.khan):

\[\frac{ 3x - 8}{4x + 20}\to \frac{3x-8}{4(x+5)}\]As there is nothing common between denominator and numerator so we cannot simply it.

OpenStudy (anonymous):

sorry I just found this.

OpenStudy (saifoo.khan):

No worries! :D

OpenStudy (anonymous):

How do I find the domain

OpenStudy (anonymous):

hi

OpenStudy (akshay_budhkar):

domain is all real numbers except for -5

OpenStudy (anonymous):

No division by zero. So what is x going to be when the denominator is zero?

OpenStudy (anonymous):

0

OpenStudy (saifoo.khan):

set the denominator equal to zero right? @Akshay_Budhkar

OpenStudy (anonymous):

no. x cannot = -5.

OpenStudy (akshay_budhkar):

Yes @sai

OpenStudy (akshay_budhkar):

@saifoo.khan

OpenStudy (anonymous):

f(x)=3x-8 over 4x+20

OpenStudy (anonymous):

@denise9423 Domain means all of the values x can take on. Here's your domain: \[x \neq-5\]

OpenStudy (akshay_budhkar):

If you have a look the function is having real values for all values except for x=5 when it is not defined as the denominator is zero

OpenStudy (anonymous):

can you show me how to get 5

OpenStudy (anonymous):

4x+20 =0 Solve for x. x=-5.

OpenStudy (anonymous):

Do you know why division by zero is not allowed?

OpenStudy (anonymous):

ok,then can i sub it also 3x-8

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

you only need to worry about the bottom.

OpenStudy (anonymous):

no to ur question ok

OpenStudy (anonymous):

Remember this: What is 1 divided by 0? In other words, what times zero can equal 1? 0*2? 0*3? There is nothing you can multiply by zero to get 1. So whenever you get these fractions and they ask you to find the domain, just think about what x has to be in order for the denominator to equal zero.

OpenStudy (anonymous):

If you have an equation in the denominator, then set that equal to zero :-)

OpenStudy (anonymous):

*and solve for x.

OpenStudy (anonymous):

like this 4x+20=0

OpenStudy (anonymous):

Yup! :-)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Or if you have 1/(x^2+2x+3) . If you have something weird like that, then set it equal to zero and use the quadratic formula to solve for your two x-values. Those x-values are what x can't equal. Sorry, I sound like I'm talking in riddles lol.

OpenStudy (anonymous):

ok so -7xyy over 28xxxxxy=0

OpenStudy (anonymous):

Are you in algebra?

OpenStudy (anonymous):

yes and Its rough

OpenStudy (anonymous):

I'm gonna be more specific with what I said, sorry my fault: Or if you have 1/(x^2+2x+3) . If you have something weird like that, then set it equal to zero and use the quadratic formula to solve for your two x-values. Those x-values are what x can't equal. Sorry, I sound like I'm talking in riddles lol. x^2+2x+3 = 0 (then use the quadratic formula and solve for x)

OpenStudy (anonymous):

That's okay!

OpenStudy (anonymous):

I don't think they'll ask you to find the domain of something with x's and y's. Or are they?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

I can help you on something like that, but first off is that denominator stuff making sense? Really sleep on that and try to remember it b/c it's super important for calculus.

OpenStudy (anonymous):

-7xyy/28xxxxxy

OpenStudy (anonymous):

Okay well what multiples to 7 and 8 have in common? What can you divide them both by?

OpenStudy (anonymous):

I mean 28.

OpenStudy (anonymous):

yes it does bc it will help me remember to put=0 so I can solve for x

OpenStudy (anonymous):

good :-)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

-7xyy/28xxxxxy You can divide the numerator and denominator by 7. Also, which has the smallest amount of x's, and then which one has the smallest amount of y's?

OpenStudy (anonymous):

3x7=28

OpenStudy (anonymous):

no 4x7=28

OpenStudy (anonymous):

The numerator has the smallest amount of x's. There is 1 of them. The denominator has the smallest amount of y's, which is 1 y. So we'll divide the numerator and denominator by 7. -1*x*y*y/4xxxxxy

OpenStudy (anonymous):

bottom has 5 x's and the top has 2 y's

OpenStudy (anonymous):

We have to cancel out the smallest amount of x's. And we also cancel out the smallest amount of y's. So we have: -1*y/4xxxx I canceled out an x on the top and bottom. Then I canceled out a y on the top and bottom.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry i have a 911 gotta chk out b back later thank you for all ur help god bless

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Anyway, -1*y/4xxxx can be re-written as: -y/4x^4 See Paul, my favorite math teacher: (I swear by his notes) http://tutorial.math.lamar.edu/Classes/Alg/IntegerExponents.aspx